OpenStudy (anonymous):

what is the inverse equation of f(x)=8sin(10x)

OpenStudy (jdoe0001):

\(\bf f(x)=8sin(10x) \implies y = 8sin(10x)\\ \quad \\ f^{-1}(x) \implies x = 8sin(10y) \implies \cfrac{x}{8} = sin(10y)\\ \implies sin^{-1}\left(\cfrac{x}{8}\right) = sin^{-1}\left(sin(10y)\right)\) I gather you can get it from there

OpenStudy (anonymous):

ok, thanks